1. If P(A) = 0.7, P(A or B) = 0.9, and P(A and B) = 0.6, then find P(B)

2. If A and B are mutually exclusive events with P(A) = 0.2 and P(B) = 0.4, find P(A or B)
Answer #1

1)

P(A or B) = P(A) + P(B) - P(A and B)

0.9 = 0.7 + P(B) - 0.6

0.9 = P(B) + 0.1

P(B) = 0.8

2)

For mutually exclusive events , P(A and B) = 0

P(A or B) = P(A) + P(B) - P(A and B)

= 0.2 + 0.4 + 0

P(A or B) = 0.6